OpenStudy (kenialove):

if m angle A = ( 4x+2 ) degrees , what is the measure of the complement of angle A ?

OpenStudy (3mar):

Complement angles are their sum of 90 degree

OpenStudy (leyb):

Complement of angle A is 90-A =90-(4x+2)

OpenStudy (leyb):

90-(4x+2) = 90-4x-2 = -4x+(90-2) = -4x+88 = -4(x-44) Hope this helped!

OpenStudy (3mar):

@kenialove

OpenStudy (kenialove):

thanks it was really helpful ! @Leyb

OpenStudy (leyb):

You're welcome! @kenialove